OpenStudy (falco276):

HELP!!!!!!!!!! segment PR has an endpoint at (25, -5) and a midpoint of (18, -1). What is the value of the x-coordinate of the other endpoint?

OpenStudy (falco276):

@Gin_Ichimaru @ganeshie8 @goformit100 @skullpatrol @DLS @phi @Preetha @canadalover_

OpenStudy (phi):

However far you go to get from the endpoint's x = 25 to get to the midpoints x=18, you go that far again to get to the other endpoint

OpenStudy (falco276):

So, if this were to be on a coordinate plane, how would I map this out?

OpenStudy (phi):

Did you get the x value of the other end point ?

OpenStudy (falco276):

No. :(

OpenStudy (phi):

you don't need to plot anything. You just have to know when you move from the endpoint to the midpoint you have to move so many steps in the x direction how many steps does it take to get from 25 to 18

OpenStudy (falco276):

7 steps?

OpenStudy (phi):

yes. (but you are getting smaller, so -7 might make more sense) now you are at the midpoint's x. the other endpoint is 7 more steps what is the x value of the other endpoint ?

OpenStudy (falco276):

wait. I'm confused. :P

OpenStudy (phi):

how many steps does it take to get from 25 to 18 7 now go another 7 steps in the same direction to get to the other endpoint

OpenStudy (falco276):

14.

OpenStudy (falco276):

Thank you. :)

OpenStudy (phi):

14 steps from 25 is the other endpoint

OpenStudy (falco276):

:?

OpenStudy (phi):

I was too quick. 14 is not the x value you want 25..... 18 ...... ? what is ? it is 7 less than 18 (or 14 less than 25)

OpenStudy (anonymous):

you may apply mid-point formula  \[(x,y) is \mid point.x=\frac{ x1+x2 }{ 2 },y=\frac{ y1+y2 }{2}\]

OpenStudy (falco276):

okay. So there is a formula required to this?

OpenStudy (anonymous):

yes

OpenStudy (phi):

if you are at the midpoint and go 7 steps up you get to 25 if you are at the midpoint and go 7 steps down you get to the other endpoint

OpenStudy (falco276):

18?

OpenStudy (phi):

18 is the midpoint

OpenStudy (falco276):

Thank you so much! :)

OpenStudy (phi):

18 is not the answer.

OpenStudy (falco276):

NOOOOOOOO!!!!!!!!!!!!!

OpenStudy (falco276):

WHY DID I DO THAT!!!!!!!!!!!

OpenStudy (falco276):

I SUBMITTED THE EXAM WITHOUT BEING CORRECT!!!!!!!!!!!!!!!!!!!!!

OpenStudy (falco276):

x.x